# Disentangled Latent Energy-Based Style Translation: An Image-Level Structural MRI Harmonization Framework

### v1

@@ Abstract
Brain magnetic resonance imaging (MRI) has been extensively employed across clinical and research fields, but often exhibits sensitivity to site effects arising from non-biological variations such as differences in field strength and scanner vendors. Numerous retrospective MRI harmonization techniques have demonstrated encouraging outcomes in reducing the site effects at image level. However, existing methods generally suffer from high computational requirements and limited generalizability, restricting their applicability to unseen MRIs. In this paper, we design a novel disentangled latent energy-based style translation (DLEST) framework for unpaired image-level MRI harmonization, consisting of (1) site-invariant image generation (SIG), (2) site-specific style translation (SST), and (3) site-specific MRI synthesis (SMS). Specifically, the SIG employs a latent autoencoder to encode MRIs into a low-dimensional latent space and reconstruct MRIs from latent codes. The SST utilizes an energy-based model to comprehend global latent distribution of a target domain and translate source latent codes toward the target domain, while SMS enables MRI synthesis with a target-specific style. By disentangling image generation and style translation in latent space, the DLEST can achieve efficient style translation. Our model was trained on T1-weighted MRIs from a public dataset (with 3,984 subjects across 58 acquisition sites/settings) and validated on an independent dataset (with 9 traveling subjects scanned in 11 sites/settings) in 4 tasks: (1) histogram and clustering comparison, (2) site classification, (3) brain tissue segmentation, and (4) site-specific MRI synthesis. Qualitative and quantitative results demonstrate the superiority of our method over several state-of-the-arts.
@@
Figure 1: The proposed disentangled latent energy-based style translation (DLEST) framework for MRI harmonization at the image level, consisting of (1) site-invariant image generation that encodes images into low-dimensional latent space and reconstruct images based on latent codes, (2) site-specific style translation that facilitates implicit style translation within the latent space, and (3) site-specific MRI synthesis that generates diverse synthetic MRIs with a given target site style.
@@ 1 Introduction
Structural magnetic resonance imaging (MRI) has found extensive application in both clinical diagnostics and research studies, primarily due to its ability to offer high spatial resolution in a non-invasive manner (Evans et al., 2020; Albayram et al., 2022; Varela et al., 2017). Large-scale data initiatives and neuroimaging studies often utilize brain MRI data pooled from various acquisition sites/settings to enhance subject cohort and boost statistical power of learning models (An et al., 2022; Tofts and Collins, 2011; Schnack et al., 2010). However, models directly trained on these multi-site datasets often encounter difficulties in achieving convergence (Glocker et al., 2019; Sun et al., 2021) and in generalizing to new data, due to the sensitivity of MRI to non-biological variations, known as site effects. The site effect can be attributed to a variety of factors, including differences in field strengths, scanner vendors, scanning protocols, and hardware upgrade and maintenance. (Glocker et al., 2019; Wachinger et al., 2021; Helmer et al., 2016). While multi-site studies, such as Alzheimer’s Disease Neuroimaging Initiative (Jack Jr et al., 2008), aim to mitigate the site effects during data acquisition through unified scanning protocols, certain variations are still inevitable, like those caused by scanner maintenance and software upgrades. Additionally, many studies rely on the retrospective pooling of multi-site data, when there was no coordination among different sites during image acquisition (Dufumier et al., 2022). One possible solution to address such a challenge is to retrospectively harmonize the imaging data as a preprocessing step (Cackowski et al., 2023; Zuo et al., 2021).
@@
Existing methods for MRI harmonization can generally be classified into two categories: (1) feature-level and (2) image-level approaches. Feature-level techniques, such as ComBat (Johnson et al., 2007) and ComBat-GAM (Pomponio et al., 2020), typically utilize empirical Bayes models and pre-extracted biological or radiomic features from a set of pre-defined regions-of-interest (ROIs). Nonetheless, the effectiveness of these methods is largely contingent on the size of datasets and the quality of feature extraction, which restricts their versatility across applications that employ diverse MRI features (Cackowski et al., 2023; An et al., 2022). On the other hand, image-level harmonization is designed for a wider range of downstream applications and is not reliant on pre-defined MRI features, which is the primary objective of this paper.
@@
For image-level harmonization, many deep generative models have been employed in recent studies (Zhu et al., 2017; Choi et al., 2020; Liu et al., 2021; Zuo et al., 2021; Cackowski et al., 2023). These methods, rather than harmonizing a set of pre-extracted MRI features, approach the harmonization problem as a pixel-to-pixel translation task at the image level. The objective is to alter the style of a source MRI to match a target MRI while preserving the original content. In this context, the style information refers to high-level features like intensity difference, contrast variations, and signal-to-noise ratio, while content information pertains to low-level image features such as anatomical structures, contours, and edges. Several state-of-the-art (SOTA) methods utilize generative adversarial networks (GANs) for image-level MRI harmonization. For instance, CycleGAN (Zhu et al., 2017) imposes cycle-consistency constraints in its loss function to enforce content preservation during unpaired style translation. StarGAN (Choi et al., 2020) employs separate encoders to learn site-specific style codes and injects them into images via adaptive instance normalization. ImUnity (Cackowski et al., 2023) incorporates additional encoder networks as extra constraints for learning latent representations to remove the site bias and retain clinical information. In addition to GAN-based models, some recent works utilize encoder-decoder models to learn disentangled latent representations for anatomical structure and contrast to perform image-level MRI harmonization (Zuo et al., 2022; Dewey et al., 2020). However, existing image-level methods usually utilize deep networks to separately learn style and content encoding in image space, leading to significantly increased training time and computational cost due to the optimization of numerous network parameters (Gulrajani et al., 2017). Furthermore, many approaches need to retrain their models entirely when applied to unseen MRI data, therefore limiting their practical generalizability (Cackowski et al., 2023).
@@
In this paper, we propose a novel disentangled latent energy-based style translation (DLEST) framework for unpaired structural MRI harmonization at the image level. As illustrated in Fig. 1, the DLEST includes (1) a site-invariant image generation module, using a latent autoencoder for encoding and generating MRIs from a lower-dimensional latent space, (2) a site-specific style translation module, using an energy-based model for implicit style translation through latent space, and (3) a site-specific MRI synthesis module for generating synthetic MRI scans with given site-specific styles. Through the disentanglement of image generation and style translation in latent space, the DLEST can achieve efficient and generalizable MRI harmonization. It can also be extended to synthesize MRIs with site-specific styles and diverse anatomical details, without extra model training.
@@
An initial version of our work has been introduced in MLMI 2023 (Wu et al., 2023). This journal paper introduces several advancements, including 1) extension of our framework for site-specific MRI synthesis, with qualitative and quantitative evaluations, 2) a thorough literature review on image-level harmonization methods and the application of energy-based models for generation tasks, 3) adding comparison with a recent state-of-the-art method, 4) experimentation with specific loss functions for content and style regularization, 5) adding comparative visualization of brain tissue segmentation results, and 6) adding quantitative comparison of computational cost across all learning methods.
@@
    Our framework can generalize to new data without model retraining since the image generation module can directly encode and reconstruct MRIs without site specification. Once trained for harmonization, our framework can natively synthesize MRIs with style of a specified site.
@@ 2.1 Image-Level MRI Harmonization
Inspired by the image-to-image translation task in the natural imaging domain, recent studies have employed the generative adversarial network (GAN) models to tackle medical data harmonization problems on the image level (Zhu et al., 2017; Choi et al., 2020; Liu et al., 2021). These methods engage the generator and discriminator networks in an adversarial game, where the generator creates synthetic images resembling the real dataset distribution, and the discriminator differentiates between synthetic and real images (Zhu et al., 2017). For instance, CycleGAN introduces a cycle-consistency constraint in its loss function for unpaired image translation and content (anatomical structure) preservation (Zhu et al., 2017). Style-encoding GAN (Liu et al., 2021), inspired by StarGAN-V2 (Choi et al., 2020), further separates the content and style encoding in the latent space, allowing the site-specific style code to be learned and injected into images via adaptive instance normalization. ImUnity (Cackowski et al., 2023) modifies the GAN structure by adding a site/scanner unlearning module to encourage the encoder to learn domain-invariant latent representations. These advancements have contributed to the continual improvement of GAN-based harmonization methods.
In addition to GAN-based models, recent studies have introduced an alternative approach that employs encoder-decoder networks to disentangle anatomical and contrast information in latent space. For instance, CALAMITI (Zuo et al., 2021) first uses T1- and T2-weighted (T1/T2-w) MRI pairs to learn global latent codes containing anatomical and contrast information, and then disentangles style and content latent codes via separate encoders and decoders. Similarly, Dewey et al.  (Dewey et al., 2020) leverage T1-w and T2-w image pairs to attain a disentangled latent space, comprising high-resolution anatomical and low-dimensional contrast components via a Randomization block. This block allows generating MRIs with identical anatomical structures but varying contrast. Zuo et al.  (Zuo et al., 2022) enhance this approach by ablating the need for paired MRI sequences. Instead, they employ 2D slices from axial and coronal views of the same MRI to provide the same contrast but different anatomical information.
These existing methods usually involve simultaneous training of multiple encoders, decoders, or sub-networks for separated style and content encoding, resulting in increased computational and time costs for model training. Besides, some methods have limited generalizability on new MRI as their models need to be retrained entirely on new datasets. Moreover, certain existing studies necessitate paired MRI scans for model training, like traveling subjects or multiple sequences, which are often unavailable in practice. In this work, we aim to enhance the efficiency of image-level harmonization and the model’s generalizability to new data, while also expanding its capacity for site-specific image synthesis.
@@ 2.2 Energy-Based Models for Image Generation
Generative models are a type of probabilistic model, which models a joint probability of the observed data and the target data. For the image generation task, a model observes image samples or their latent representations and learns to capture the true data distribution to generate new images that resemble the observed samples (Pang et al., 2020; Du and Mordatch, 2019). Probabilistic models employ statistical inference methods, such as maximum likelihood estimation (MLE), to estimate parameters that maximize the likelihood of the observed data under an assumed distribution (Ng and Jordan, 2001). However, these models need to be correctly normalized. This process can sometimes involve the evaluation of intractable integrals over all possible variable configurations, which can be computationally challenging. Therefore, LeCun et al.(LeCun et al., 2006) use energy-based learning as an alternative framework for both discriminative and generative tasks. Energy-based models (EBMs) capture data dependencies by associating a scalar energy value (a measure of compatibility) to each configuration of variables (LeCun et al., 2006). Unlike probabilistic models, EBMs do not require normalization. For generative tasks, the training phase of EBMs involves learning an appropriate energy function that assigns low energy values to observed data within the real data distribution and high energy values otherwise, during which data is typically sampled via Markov Chain Monte Carlo (MCMC) sampling methods. During inference, the EBMs generate data by exploring the energy surface to sample configurations that minimize the energy function.
@@
Over the years, many studies have employed EBMs for image generation and improved on various aspects. Du et al.(Du and Mordatch, 2019) propose to utilize Langevin Dynamics, a gradient-based MCMC method, for more efficient sampling and mixing on high-dimensional image domains. Pang et al.(Pang et al., 2020) further improve the sampling efficiency and the MCMC mixing by combining an EBM with a top-down network (decoder) in the low-dimensional latent space. The EBM serves to refine the simple isotropic Gaussian prior to be closer to the data distribution, leading to improved quality during image generation. Xiao et al.(Xiao et al., 2020) propose a symbiosis of Variational Autoencoder (VAE) and EBM to leverage the advantages of both. The VAE captures the overall model structure of the data distribution and provides an improved latent initialization of the MCMC sampling, which is used in EBM. The EBM, on the other hand, can exclude non-data-like samples and refine the latent representation for VAE to decode. A recent work (Zhao and Chen, 2021) utilizes this symbiotic combination of EBM and autoencoder for the task of male-to-female face translation. In this context, the EBM learns the latent distribution of a target image domain and translates the source latent representation towards this target distribution. The underlying assumption is that two image domains share certain common latent representations. The EBM is then able to implicitly translate latent codes that encapsulate domain differences while preserving shared latent codes that signify domain similarities. Inspired by the success of EBMs in natural image-to-image translation, our method adopts a similar assumption. We posit that MRIs from different domains share some common latent representations that convey content (anatomical) information. We propose training an EBM to translate the site-specific style information from one domain to another within a low-dimensional latent space.
@@ 3.1.1 Datasets
Our study utilizes two public datasets, including (1) Open Big Healthy Brains (OpenBHB) (Dufumier et al., 2022) and (2) Strategic Research Program for Brain Science (SRPBS) (Tanaka et al., 2021). Specifically, the OpenBHB, a large-scale brain MRI dataset, comprises 3,984 T1-weighted MRIs from healthy subjects acquired from over 58 centers globally. The SRPBS dataset contains 108 T1-weighted structural MRIs acquired from 9 healthy traveling subjects across 12 scanners/settings. In this work, MR scans of 9 subjects from 11 acquisition sites/settings in SRPBS are used, while one duplicated site (YC2) is removed. Further details can be found ${}^{,}$ [^2] https://www.synapse.org/#!Synapse:syn22317076/wiki/605026.
@@ 3.1.2 Data Preprocessing
For these two datasets, each MRI scan undergoes minimal preprocessing using FSL ANAT pipeline (Smith et al., 2004), which encompasses bias field correction, brain extraction, and 9 degrees of freedom linear registration to the MNI-152 template with $1\,mm$ isotropic resolution. Brain tissue segmentation is performed on SRPBS images using FSL FAST pipeline, with each brain segmented into white matter (WM), gray matter (GM), and cerebrospinal fluid (CSF). For each MRI from OpenBHB, we selected 10 slices in axial view from the center of each MRI. For SRPBS with limited MRIs, we chose 15 slices from each volume. Additionally, all these 2D slices were zero-padded to have the size of $256\times 256$.

### v2

@@ Abstract
Brain magnetic resonance imaging (MRI) has been extensively employed across clinical and research fields, but often exhibits sensitivity to site effects arising from non-biological variations such as differences in field strength and scanner vendors. Numerous retrospective MRI harmonization techniques have demonstrated encouraging outcomes in reducing the site effects at image level. However, existing methods generally suffer from high computational requirements and limited generalizability, restricting their applicability to unseen MRIs. In this paper, we design a novel disentangled latent energy-based style translation (DLEST) framework for unpaired image-level MRI harmonization, consisting of (a) site-invariant image generation (SIG), (b) site-specific style translation (SST), and (c) site-specific MRI synthesis (SMS). Specifically, the SIG employs a latent autoencoder to encode MRIs into a low-dimensional latent space and reconstruct MRIs from latent codes. The SST utilizes an energy-based model to comprehend global latent distribution of a target domain and translate source latent codes toward the target domain, while SMS enables MRI synthesis with a target-specific style. By disentangling image generation and style translation in latent space, the DLEST can achieve efficient style translation. Our model was trained on T1-weighted MRIs from a public dataset (with 3,984 subjects across 58 acquisition sites/settings) and validated on an independent dataset (with 9 traveling subjects scanned in 11 sites/settings) in four tasks: histogram and feature visualization, site classification, brain tissue segmentation, and site-specific structural MRI synthesis. Qualitative and quantitative results demonstrate the superiority of our method over several state-of-the-arts.
@@
Figure 1: The proposed disentangled latent energy-based style translation (DLEST) framework for MRI harmonization at the image level, consisting of (a) site-invariant image generation that encodes images into low-dimensional latent space and reconstruct images based on latent codes, (b) site-specific style translation that facilitates implicit style translation within the latent space, and (c) site-specific MRI synthesis that generates diverse synthetic MRIs with a given target site style.
@@ 1 Introduction
Structural magnetic resonance imaging (MRI) has found extensive application in both clinical diagnostics and research studies, primarily due to its ability to offer high spatial resolution in a non-invasive manner [10, 1, 41]. Large-scale data initiatives and neuroimaging studies often utilize brain MRI data pooled from various acquisition sites/settings to enhance subject cohort and boost the statistical power of learning models [2, 40, 33]. However, models directly trained on these multi-site datasets often encounter difficulties in achieving convergence [12, 38] and in generalizing to new data, due to the sensitivity of MRI to non-biological variations, known as site effects. The site effect can be attributed to a variety of factors, including differences in field strengths, scanner vendors, scanning protocols, and hardware upgrade and maintenance. [12, 42, 16]. While multi-site studies, such as Alzheimer’s Disease Neuroimaging Initiative [17], aim to mitigate site effects during data acquisition through unified scanning protocols, certain variations are inevitable, like those caused by scanner maintenance or software upgrades. Additionally, many studies rely on the retrospective pooling of multi-site data, when there was no coordination among different sites during image acquisition [9]. One possible solution to address such a challenge is to retrospectively harmonize the imaging data as a preprocessing step [3, 49].
@@
Existing methods for MRI harmonization can generally be classified into two categories: feature-level approaches and image-level approaches. Feature-level techniques, such as ComBat [18] and ComBat-GAM [30], typically utilize empirical Bayes models and pre-extracted biological or radiomic features from a set of pre-defined regions-of-interest (ROIs). Nonetheless, the effectiveness of these methods is largely contingent on the size of datasets and the quality of feature extraction, which restricts their versatility across applications that employ diverse MRI features [3, 2]. On the other hand, image-level harmonization is designed for a wider range of downstream applications. It is not reliant on pre-defined MRI features, which is the primary objective of this paper.
@@
For image-level harmonization, many deep generative models have been employed in recent studies [48, 5, 23, 49, 3]. These methods, rather than harmonizing a set of pre-extracted MRI features, approach the harmonization problem as a pixel-to-pixel translation task at the image level. The objective is to alter the style of a source MRI to match a target MRI while preserving the original content. In this context, the style information refers to high-level features like intensity difference, contrast variations, and signal-to-noise ratio, while content information pertains to low-level image features such as anatomical structures, contours, and edges. Several state-of-the-art (SOTA) methods utilize generative adversarial networks (GANs) for image-level MRI harmonization. For instance, CycleGAN [48] imposes cycle-consistency constraints in its loss function to enforce content preservation during unpaired style translation. StarGAN [5] employs separate encoders to learn site-specific style codes and injects them into images via adaptive instance normalization. ImUnity [3] incorporates additional encoder networks as extra constraints for learning latent representations to remove the site bias and retain clinical information. In addition to GAN-based models, some recent works utilize encoder-decoder models to learn disentangled latent representations for anatomical structure and contrast to perform image-level MRI harmonization [50, 7]. However, existing image-level methods usually utilize deep networks to separately learn style and content encoding in image space, leading to significantly increased training time and computational cost due to the optimization of numerous network parameters [14]. Furthermore, many approaches need to retrain their models entirely when applied to unseen MRI data, therefore limiting their practical generalizability [3].
@@
In this paper, we propose a novel disentangled latent energy-based style translation (DLEST) framework for unpaired structural MRI harmonization at the image level. As illustrated in Fig. 1, the DLEST includes (a) a site-invariant image generation module, using a latent autoencoder for encoding and generating MRIs from a lower-dimensional latent space, (b) a site-specific style translation module, using an energy-based model for implicit style translation through latent space, and (c) a site-specific MRI synthesis module for generating synthetic MRIs with given site-specific styles. Through disentangling image generation and style translation in latent space, the DLEST can achieve efficient and generalizable MRI harmonization. It can also be extended to synthesize MRIs with site-specific styles and diverse anatomical details, without extra model training.
@@
An initial version of our work has been introduced in MLMI 2023 [45]. This journal paper introduces several advancements, including 1) extension of our framework for site-specific MRI synthesis, with qualitative and quantitative evaluations, 2) a thorough literature review on image-level harmonization methods and the application of energy-based models for generation tasks, 3) adding comparison with a recent state-of-the-art method, 4) experimentation with specific loss functions for content and style regularization, 5) adding comparative visualization of brain tissue segmentation results, 6) including experimental analysis on the influence of target site selection, and 7) adding quantitative comparison of computational cost across all learning methods.
@@
    Our framework can generalize to new data without model retraining since the image generation module can directly encode and reconstruct MRIs without site specification, while the style translation module can be fine-tuned during the testing stage with minimal computational cost. Once trained for harmonization, our method can synthesize MRIs with the style of a specified site.
@@ 2.1 Image-Level MRI Harmonization
Inspired by the image-to-image translation task in the natural imaging domain, recent studies have employed the generative adversarial network (GAN) models to tackle medical data harmonization problems on the image level [48, 5, 23]. These methods engage the generator and discriminator networks in an adversarial game, where the generator creates synthetic images resembling the real dataset distribution, and the discriminator differentiates between synthetic and real images [48]. For instance, CycleGAN introduces a cycle-consistency constraint in its loss function for unpaired image translation and content (anatomical structure) preservation [48]. Style-encoding GAN [23], inspired by StarGAN-V2 [5], further separates the content and style encoding in the latent space, allowing the site-specific style code to be learned and injected into images via adaptive instance normalization. ImUnity [3] modifies the GAN structure by adding a site/scanner unlearning module to encourage the encoder to learn domain-invariant latent representations. These advancements have contributed to the continual improvement of GAN-based harmonization methods.
In addition to GAN-based models, recent studies have introduced an alternative approach that employs encoder-decoder networks to disentangle anatomical and contrast information in latent space. For instance, CALAMITI [49] first uses T1- and T2-weighted (T1/T2-w) MRI pairs to learn global latent codes containing anatomical and contrast information, and then disentangles style and content latent codes via separate encoders and decoders. Similarly, Dewey et al.  [7] leverage T1-w and T2-w image pairs to attain a disentangled latent space, comprising high-resolution anatomical and low-dimensional contrast components via a Randomization block. This block allows generating MRIs with identical anatomical structures but varying contrast. Zuo et al.  [50] enhance this approach by ablating the need for paired MRI sequences. Instead, they employ 2D slices from axial and coronal views of the same MRI to provide the same contrast but different anatomical information.
These existing methods usually involve simultaneous training of multiple encoders, decoders, or sub-networks for separated style and content encoding, resulting in increased computational and time costs for model training. Besides, some methods have limited generalizability on new MRI as their models need to be retrained entirely on new datasets. Moreover, certain existing studies necessitate paired MRI scans for model training, like traveling subjects or multiple MRI sequences, which are often unavailable in practice. In this work, we aim to enhance the efficiency of image-level harmonization and the model’s generalizability to new data, while also expanding its capacity for site-specific image synthesis.
@@ 2.2 Energy-Based Models for Image Generation
Generative models are a category of probabilistic models that represent the joint probability distribution of both observed and target data. For the image generation task, a model observes image samples or their latent representations and learns to capture the true data distribution to generate new images that resemble the observed samples [27, 8]. Probabilistic models employ statistical inference methods, such as maximum likelihood estimation (MLE), to estimate parameters that maximize the likelihood of the observed data under an assumed distribution [26]. However, these models need to be correctly normalized. This process can sometimes involve the evaluation of intractable integrals over all possible variable configurations, which can be computationally challenging. Therefore, LeCun et al.[21] use energy-based learning as an alternative framework for both discriminative and generative tasks. Energy-based models (EBMs) capture data dependencies by associating a scalar energy value (a measure of compatibility) to each configuration of variables [21]. Unlike probabilistic models, EBMs do not require normalization. For generative tasks, the training phase of EBMs involves learning an appropriate energy function that assigns low energy values to observed data within the real data distribution and high energy values otherwise, during which data is typically sampled via Markov Chain Monte Carlo (MCMC) sampling methods. During inference, the EBMs generate data by exploring the energy surface to sample configurations that minimize the energy function.
@@
Over the years, many studies have employed EBMs for image generation and improved on various aspects. Du et al.[8] propose to utilize Langevin Dynamics, a gradient-based MCMC method, for more efficient sampling and mixing on high-dimensional image domains. Pang et al.[27] further improves the sampling efficiency and the MCMC mixing by combining an EBM with a top-down network (decoder) in the low-dimensional latent space. The EBM serves to refine the simple isotropic Gaussian prior to be closer to the data distribution, leading to improved quality during image generation. Xiao et al.[46] propose a symbiosis of Variational Autoencoder (VAE) and EBM to leverage the advantages of both. The VAE captures the overall model structure of the data distribution and provides an improved latent initialization of the MCMC sampling, which is used in EBM. The EBM, on the other hand, can exclude non-data-like samples and refine the latent representation for VAE to decode. A recent work [47] utilizes this symbiotic combination of EBM and autoencoder for the task of male-to-female face translation. In this context, the EBM learns the latent distribution of a target image domain and translates the source latent representation towards this target distribution. The underlying assumption is that two image domains share certain common latent representations. The EBM is then able to implicitly translate latent codes that encapsulate domain differences while preserving shared latent codes that signify domain similarities. Inspired by the success of EBMs in natural image-to-image translation, our method adopts a similar assumption. We posit that MRIs from different domains share some common latent representations that convey content (anatomical) information. We propose training an EBM to translate the site-specific style information from one domain to another within a low-dimensional latent space.
@@ 3.1.1 Datasets
Our study utilizes two public datasets, including (1) Open Big Healthy Brains (OpenBHB) [9] and (2) Strategic Research Program for Brain Science (SRPBS) [39]. Specifically, the OpenBHB, a large-scale brain MRI dataset, comprises 3,984 T1-weighted MRIs from healthy subjects acquired from over 58 centers globally. The SRPBS dataset contains 108 T1-weighted structural MRIs acquired from 9 healthy traveling subjects across 12 scanners/settings. In this work, MR scans of 9 subjects from 11 acquisition sites/settings in SRPBS are used, while one duplicated site (YC2) is removed. Further details can be found online [^1] https://baobablab.github.io/bhb/, [^2] https://www.synapse.org/#!Synapse:syn22317076/wiki/605026.
@@ 3.1.2 Data Preprocessing
For these two datasets, each MRI scan undergoes minimal preprocessing using FSL ANAT pipeline [36], which encompasses bias field correction, brain extraction, and 9 degrees of freedom linear registration to the MNI-152 template with $1\,mm$ isotropic resolution. Brain tissue segmentation is performed on SRPBS images using FSL FAST pipeline, with each brain segmented into white matter (WM), gray matter (GM), and cerebrospinal fluid (CSF). For each MRI from OpenBHB, we select 10 slices in axial view from the center of each MRI. We utilize the original training and test data split specified in [9], which eventually results in $32,270$ 2D MRI slices for training and $7,570$ MRI slices for validation. For SRPBS with limited traveling subject MRIs, we choose 14 slices from each volume, resulting in $1,386$ 2D MRI slices. Additionally, all these 2D slices are zero-padded to have the size of $256\times 256$.